OpenStudy (anonymous):

Let h=v0^2/4.9 sin(theta) cos(theta) model the horizontal distance in meters traveled by a projectile. If the initial velocity is 44 meters/second, which equation would you use to find the angle needed to travel 150 meters? A. 8.98sin(2 theta) = 150 B. 395.10sin(2 theta) = 150 C. 150sin(2 theta) = 150 D. 197.55sin(2 theta) = 150 @terenzreignz what do u think? this one, i don't get.. like at all :/ haha :P

OpenStudy (anonymous):

i'm kinda leaning towards B... but i'm not too sure on this whole process :/

terenzreignz (terenzreignz):

I see... hang on a sec.

OpenStudy (anonymous):

kk :)

terenzreignz (terenzreignz):

Okay... you clearly see that v0 = 44, right?

OpenStudy (anonymous):

yes :) do i need to square that? :/

OpenStudy (anonymous):

like 44^2 ? :/ not sure if that's what I'm supposed to do?

terenzreignz (terenzreignz):

Okay, tell you what, let's put this range formula in its more traditional form... \[\large \frac{v_o^2}{4.9}\sin(\theta)\cos(\theta)\]

OpenStudy (anonymous):

kk :)

terenzreignz (terenzreignz):

So, we can multiply and divide by 2... to keep things the same \[\large \frac{v_o^2}{4.9}\cdot\frac12\cdot2\sin(\theta)\cos(\theta)\]

terenzreignz (terenzreignz):

Giving us \[\large \frac{v_o^2}{9.8}\cdot2\sin(\theta)\cos(\theta)\]

OpenStudy (anonymous):

seriously, latex rocks! :) haha oaky! :D

terenzreignz (terenzreignz):

\[\large \frac{v_o^2}{9.8}\cdot\color{red}{2\sin(\theta)\cos(\theta)}\] This bit here is the very definition of \(\large \sin(2\theta)\) So we get \[\large \frac{v_o^2}{9.8}\cdot\sin(2\theta)\] Now plug in your \(\large v_o = 44 \)

OpenStudy (anonymous):

okay... ermm sooo 44^2/9.8 x sin(2 theta) ?

terenzreignz (terenzreignz):

yup...

OpenStudy (anonymous):

so 1936/9.8=197.55 right?

terenzreignz (terenzreignz):

yes.

OpenStudy (anonymous):

so the answer is D. 197.55sin(2 theta) = 150 ?? :D

terenzreignz (terenzreignz):

What does your heart tell you?

OpenStudy (anonymous):

that it's D :) hehe

OpenStudy (anonymous):

hehe thank youu :)